Can you please solve the system of 4x-2y=9 and y=2x+5 ? Can you also please solve the system 4x-y=8 and 2x-4=3y ? Thanks
Answer by oscargut(2103)

4x-2y=9
y=2x+5
Answer: No solution

4x-y=8 2x-
2x-4=3y
answer: (2,0)
I can help you with more questions at: mthman@gmail.com
Thanks